COMPLETE LIST OF PUBLICATIONS

Krzysztof Galicki

• Articles published or accepted for publication

1. On Sasakian η-Einstein Sasakian Geometry, Communications in Mathematical Physics 262 (2006), 177-208; (with C. P. Boyer and P. Matzeu).

2. Sasakian geometry Einstein metrics on spheres, proceedings of the Short Program on Riemannian Geometry, Centre de Recherche Mathématiques, Université de Montréal, June 28-July 16, 2004; to appear; (with C. P. Boyer).

3. Einstein metrics on spheres, Annals of Mathematics (2005), to appear (with C. P. Boyer and J. Kollár).

4. Einstein metrics on exotic spheres in dimension 7,11, and 15, Experimental Mathematics 14 (2005), 61-66; (with C. P. Boyer, J. Kollár, and E. Thomas).

5. Sasakian Geometry, Hypersurface Singularities, and Einstein Metrics, Rendiconti del Circolo Matematico di Palermo (2), Suppl. No. 75 (2005), 57-87; (with C. P. Boyer).

6. Toric self-dual Einstein metrics as quotients, Communications in Mathematical Physics 253 (2005), 337-370; (with C. P. Boyer, D. M. J. Calderbank, and P. Piccinni).

7. On Positive Sasakian Geometry, Geometriae Dedicata 101 (2003), 93-102; (with C. P. Boyer and M. Nakamaye).

8. New Einstein Metrics on 8#(S^{2} × S^{3}), Differential Geometry and Its Applications 19 (2003); 245-251; (with C. P. Boyer).

9. The Geometry of Sasakian-Einstein 5-Manifolds, Mathematische Annalen 325 (2003), 485-524; (with C. P. Boyer and M. Nakamaye).

10. Sasakian Geometry, Homotopy Spheres, and Positive Ricci Curvature, Topology 42 (2003), 981-1002; (with C. P. Boyer and M. Nakamaye).

11. Einstein Metrics on Rational Homology 7-Spheres, Annales de l’Institut Fourier 52 (2002), 1569-1584; (with C. P. Boyer and M. Nakamaye).

12. Rational Homology 5-Spheres with Positive Ricci Curvature, Mathematical Research Letters 9 (2002), 521-528; (with C. P. Boyer).

13. Sasakian-Einstein Structures on #9(S^{2} × S^{3}), Transactions of the American Mathematical Society 354 (2002), 2983-2996; (with C. P. Boyer and M. Nakamaye).

14. 3-Sasakian Geometry, Nilpotent Orbits, and Exceptional Quotients, Annales of Global Analysis and Geometry 21 (2002), 85-110; (with C. P. Boyer and P. Piccinni).

15. New Einstein Metrics in Dimension Five, Journal of Differential Geometry 57 (2001); 443-463; (with C. P. Boyer).

16. Contact Geometry and Einstein Manifolds, Proceedings of the American Mathematical Society 129 (2001), 2419-2430; (with C. P. Boyer).

17. On Sasakian-Einstein Geometry, International Journal of Mathematics 11 (2000), 873-909; (with C. P. Boyer).

18. A Note on Contact Toric Geometry, Journal of Geometry and Physics 35 (2000), 288-298; (with C. P. Boyer).

19. 3-Sasakian Manifolds, Encyclopaedia of Mathematics, Supplement II, pp. 449-451, Kluwer Academic Publishers, (2000).

20. 3-Sasakian Manifolds; Surveys in Differential Geometry VI: Essays on Einstein Manifolds A supplement to the Journal of Differential Geometry, pp. 123-184, (eds. C. LeBrun, M. Wang); International Press, Cambridge (1999); (with C. P. Boyer).

21. Quaternionic geometry and 3-Sasakian manifolds, Quaternionic structures in mathematics and physics (Trieste, 1994), 7–24 (electronic), Int. Sch. Adv. Stud. (SISSA), Trieste, (1998); (with C. P. Boyer and B. Mann).

22. Hypercomplex Structures From 3-Sasakian Structures, Journal für die reine und angewandte Mathematik 501 (1998), 115-141; (with C. P. Boyer and B. Mann).

23. A Note on Smooth Toral Reductions of Spheres, Manuscripta Mathematica 95 (1998), 149-158. (with C. P. Boyer and B. Mann).

24. Compact 3-Sasakian 7-Manifolds with Arbitrary Second Betti Number, Inventiones Mathematicae 131 (1998), 321-344; (with C. P. Boyer, B. Mann, and E. Rees).

25. The Twistor Space of a 3-Sasakian Manifolds, International Journal of Mathematics 8 (1997), 31-60; (with C. P. Boyer).

26. Einstein Manifolds of Positive Scalar Curvature with Arbitrary Second Betti Number, Balkan Journal of Geometry and Its Applications 1, no. 2 (1996), 1-8. (with C. P. Boyer, B. Mann, and E. Rees).

27. On Betti Numbers of 3-Sasakian Manifolds, Geometriae Dedicata 63 (1996), 45-68; (with S. Salamon).

28. On Strongly Inhomogeneous Einstein Manifolds, Bulletin of London Mathematical Society 28 (1996), 401-408; (with C. P. Boyer and B. Mann).

29. Hypercomplex Structures on Stiefel Manifolds, Annals of Global Analysis and Geometry 14 (1996), 81-105; (with C. P. Boyer and B. Mann).

30. Some New Examples of Inhomogeneous Hypercomplex Manifolds, Mathematical Research Letters 1 (1994), 531-538; (with C. P. Boyer and B. Mann).

31. Geometry and topology of 3-Sasakian manifolds, Journal für die reine und angewandte Mathematik 455 (1994), 183-220; (with C. P. Boyer and B. Mann).

32. New Examples of Inhomogeneous Einstein Manifolds of Positive Scalar Curvature, Mathematical Research Letters 1 (1994), 115-121; (with C. P. Boyer and B. Mann).

33. 3-Sasakian Manifolds, Proceedings of Japan Academy vol. 69, Ser. A, (1993), 335-340; (with C. P. Boyer and B. Mann).

34. Quaternionic Quotients and Einstein Manifolds, in Communications in Analysis and Geometry vol. 1, no. 2, (1993), 1-51; (with C. P. Boyer and B. Mann).

35. Nonzero Scalar Curvature Generalizations of the ALE Instantons, Journal of Mathematical Physics 33 (1992), 1765-1771; (with T. Nitta).

36. Geometry of the Scalar Couplings in N = 2 Supergravity Models, Classical and Quantum Gravity 9 (1992), 27-40.

37. Multicenter Metrics with Negative Cosmological Constant, Classical and Quantum Gravity 8 (1991), 1529-1543.

38. Duality and Yang-Mills Fields on quaternionic Kähler manifolds, Journal of Mathematical Physics 32 (1991), 1263-1268; (with Y. S. Poon).

39. Quaternionic Reduction and Quaternionic Orbifolds, Mathematische Annalen 282 (1988), 1-21; (with B.H. Lawson, Jr.)

40. A Generalization of Momentum Mapping Construction for Quaternionic Kähler Manifolds, Communications in Mathematical Physics 108 (1987), 117-138.

41. New Matter Couplings in N=2 Supergravity, Nuclear Physics B289 (1987), 573-588.

42. Hyperkähler and Quaternionic Kähler Nonlinear Sigma Models, Nuclear Physics B271 (1986), 402-416.

43. CP (n) nonlinear sigma models with holomorphic constraints in d=2,4 dimensions, Physical Reviews D29 (1984), 1809-1818.

• Articles submitted for publication

1. Einstein metric on rational homology spheres, submitted to the Journal of Differential Geometry; (with C. P. Boyer).

2. Highly connected manifolds with positive Ricci curvature, submitted to Geometry and Topology; (with C. P. Boyer).

• To be submitted and in preparation

1. Constructions in Sasakian Geometry, in preparation; (with C. P. Boyer and L. Ornea).

2. Sasakian Manifolds, research monograph for the Oxford University Press; in preparation (with C. P. Boyer).

3. On Sasakian Futaki Invariants, in preparation; (with C. P. Boyer and S. Simanca).

4. Sasakian Geometry in Five Dimensions, in preparation; (with C. P. Boyer).

Last updated: February 1, 2006